OpenStudy (anonymous):

A square wall, is covered with square tiles. There are 85 tiles altogether along the two diagonals. How many tiles are there on the whole wall???

OpenStudy (anonymous):

the tiles are of equal sizes.

OpenStudy (shane_b):

There would be 85 tiles altogether if you summed up the diagonals and didn't count the center tile twice. At any rate the sides must be 86/2 tiles wide and the area (number of tiles) will be \[A_{square}=s^2=(\frac{86}{2})^2\]

OpenStudy (shane_b):

I had to think about it too :)

OpenStudy (anonymous):

what does the 's' and 'A' STAND for?

OpenStudy (shane_b):

A=area, s= side

OpenStudy (anonymous):

thank u .... appreciate it :)

OpenStudy (anonymous):

There are 43 tiles along a diagonal? Then surely the area is\[(\frac{1}{\sqrt{2}}s)^2\] As you, i probably didn't understand the question.

OpenStudy (anonymous):

With s=43=diagonal

OpenStudy (anonymous):

@henpen can u explain your formula?

OpenStudy (anonymous):

And use